# Comparative Insights on Adversarial Machine Learning from Industry and Academia: A User-Study Approach

### v1

@@ IV-B2 Ethical Considerations
Our data collection process for our study was organized in such a way that ethical considerations were taken into account at every step of the way. Our survey script, acknowledgments, posters for the study, and other relevant materials went through an approval process by the Carnegie Mellon University Institutional Review Board (IRB). The participants were compensated 50 USD for the one hour in-person survey. No personally identifiable information and audio/video was recorded during this process.
@@ IV-C2 Participant’s View: CTF Insights
TABLE III: CTF Platform: Concepts vs. Count or participants (n out of 15)
@@
For statistical analysis, we used Fisher’s Exact test to assess the correlation between the participant’s background and their preference of CTF as means of learning AML, which is our first hypothesis. For this, we selected Q2 and Q28 from the survey (see Appendix A-D) where Q2 represents the educational background of the participants and Q28 represents their preference for CTF. After converting the 5-point Likert responses (1: strongly disagree, 2: disagree, 3:neutral, 4:agree and 5: strongly agree) to a 3-category response (Yes, No and Neutral) for a clearer view, Fishers Exact Test (see Appendix A-D) yielded an $\alpha$ value of $0.006$. Since the value of $\alpha$ is less than $0.05$, we find evidence consistent with H1:
@@
For H2, we analyzed the responses to Q7 and Q28 (see Appendix A-D), where Q7 represents the’ background of the CTFs of the participants and Q28 represents their preference for CTFs. Using Fisher’s Exact Test, we obtained $\alpha=0.6$, which indicates that we did not find evidence supporting an association between prior CTF usage and preference for CTFs as an educational tool in this sample for H2:
@@ VI Future Work
As we saw in Sections III and IV, current tools for AML education are insufficient for professionals and students, resulting in a need for some resources that provide hands-on training. We proposed a novel Capture-the-Flag (CTF) educational model guidelines tailored for AML education while incorporating insights from our study. Considering the insights from our study, we suggest the following revisions to our existing CTF implementation for the future:
@@ A-B Questionnaire for second study
This is a $60$ minutes user study as part of a project at an IRB-approved academic research study related to the a widely used educational CTF platform platform. As part of this survey, we would be collecting data based on your answers to the following questions but it will not contain any personally identifiable information that may link back to you. We will also not collect any type of other information such as audio, video, keystrokes, mouse movement, etc. that may be concerning to many people’s privacy. If you have questions, please contact the corresponding author listed on the first page.
@@ A-D1 First Study
Statistical analysis of correlation between experience in cybersecurity and participant’s belief regarding relevance of cybersecurity in ML.Here, F1=Extremely relevant, F2=Moderately relevant, F3=Relevant, F4=Not Relevant.

### v2

@@ IV-B2 Ethical Considerations
Our data collection process for our study was organized in such a way that ethical considerations were taken into account at every step of the way. Our survey script, acknowledgments, posters for the study, and other relevant materials went through an approval process by the Carnegie Mellon University Institutional Review Board (IRB). The participants were compensated 50 USD for the one-hour in-person survey. No personally identifiable information or audio/video was recorded.
@@ IV-C2 Participant’s View: CTF Insights
TABLE III: CTF Platform: Concepts vs. Count of participants (n=15)
@@
For statistical analysis, we used Fisher’s exact test to assess the correlation between the participant’s background and their preference for CTF as a means of learning AML, which is our first hypothesis. For this, we selected Q2 and Q28 from the survey (see Appendix A-D) where Q2 represents the educational background of the participants and Q28 represents their preference for CTF. After converting the 5-point Likert responses (1: strongly disagree, 2: disagree, 3:neutral, 4:agree and 5: strongly agree) to a 3-category response (Yes, No and Neutral) for a clearer view, Fishers Exact Test (see Appendix A-D) yielded an $\alpha$ value of $0.006$. Since the value of $\alpha$ is less than $0.05$, we find evidence consistent with H1:
@@
For H2, we analyzed the responses to Q7 and Q28 (see Appendix A-D), where Q7 represents the background of the CTF of the participants and Q28 represents their preference for CTFs. Using Fisher’s Exact Test, we obtained $\alpha=0.6$, which indicates that we did not find evidence supporting an association between prior CTF usage and preference for CTFs as an educational tool in this sample for H2:
@@ VI Future Work
As we saw in Sections III and IV, current tools for AML education are insufficient for professionals and students, resulting in a need for some resources that provide hands-on training. We proposed a novel CTF educational model tailored for AML education while incorporating insights from our study. Considering the insights from our study, we suggest the following revisions to our existing CTF implementation for the future:
@@ A-B Questionnaire for second study
This is a $60$-minute user study as part of a project at an IRB-approved academic research study related to a widely used educational CTF platform. As part of this survey, we would be collecting data based on your answers to the following questions but it will not contain any personally identifiable information that may link back to you. We will also not collect any type of other information such as audio, video, keystrokes, mouse movement, etc. that may be concerning to many people’s privacy. If you have questions, please contact the corresponding author listed on the first page.
@@ A-D1 First Study
Statistical analysis of correlation between experience in cybersecurity and participants’ belief regarding relevance of cybersecurity in ML. Here, F1=Extremely relevant, F2=Moderately relevant, F3=Relevant, F4=Not Relevant.